1.) write an equation of the line that passes through (4, 6) and is parallel to the line that passes through (6, -6) AND (10, -4)

2.)write an equation of the line that passes through (6, 10) and is parallel to the line that passes through (4, -6) AND (3, -4)

Answer by checkley75(3666)

FIRST WE NEED TO FIND TRHE SLOPE OF THE LINE THROUGH (6,-6) & (10,4)
(4+6)/(10-6)=10/4=2.5
ANY PARALLEL LINE WILL ALSO HAVE A SLOPE=2.5.
NOW WE USE THE X,Y VALUESIN THE LINE EQUATION
Y=mX+b WHERE m=SLOPE & SOLVE FOR THE Y INTERCEPT(b).
6=2.5*4+b
6=10+b
6-10=b
-4=b
THUS THIS EQUATION IS:
Y=2.5X-4